OpenStudy (anonymous):

Solve for the indicated variable. A = 1 + prt; for r

OpenStudy (anonymous):

A - 1 = prt (A -1) / pt = r

OpenStudy (anonymous):

The goal is to keep your eye on the prize, in this case the prize is r. First try to get it by itself by going backwards through PEMDAS (order of operations) Subtract the 1 over to the other side \[A-1=prt\] Now divide the p and t to the other side \[(A-1)/(pt) = r\] And you're done.